OpenStudy (anonymous):

what would be the pressure of a given mass of a gas, if its volume and temperature are doubled ............. ?

OpenStudy (joannablackwelder):

P1V1/T1=P2V2/T2 when mass is constant.

OpenStudy (joannablackwelder):

What happens when we sub in 2V1 for V2

OpenStudy (joannablackwelder):

And 2T1 for T2?

OpenStudy (anonymous):

He wants to complement the question one word

OpenStudy (joannablackwelder):

Ok, we are getting there.

OpenStudy (anonymous):

so the answer : mass is constant.?

OpenStudy (joannablackwelder):

No, that is an assumption, since we aren't told how it will change.

OpenStudy (joannablackwelder):

Do you see how when the volume is doubled, that would be 2 times the original volume?

OpenStudy (anonymous):

yes

OpenStudy (joannablackwelder):

Ok, cool. And the same for temp?

OpenStudy (anonymous):

yup

OpenStudy (joannablackwelder):

Ok, so, can you plug those into the equation?

OpenStudy (anonymous):

P1V1/T1=P2V2/T2

OpenStudy (joannablackwelder):

That is the original equation. Can you plug in the other expressions we came up with?

OpenStudy (anonymous):

I don't understand sorry

OpenStudy (joannablackwelder):

It's cool. So, since V2 can be represented by 2V1, P1V1/T1=P2(2V1)/T2

OpenStudy (joannablackwelder):

Does that step make sense?

OpenStudy (anonymous):

yes

OpenStudy (joannablackwelder):

:-) Can you try plugging in the expression for T2?

OpenStudy (anonymous):

T2= P1V2T1/P2V1

OpenStudy (joannablackwelder):

Not solve, substitute.

OpenStudy (melodious):

joanna

OpenStudy (anonymous):

:c

OpenStudy (melodious):

that question i didnt understand

OpenStudy (melodious):

:c

OpenStudy (anonymous):

He wants to complement the question one word

OpenStudy (melodious):

ur a boy?

OpenStudy (anonymous):

yup

OpenStudy (joannablackwelder):

P1V1/T1=P2(2V1)/(2T1) Do you see how I subbed that in?

OpenStudy (joannablackwelder):

Since T2 is 2T1

OpenStudy (anonymous):

oh, yeah

OpenStudy (joannablackwelder):

:-) So, can you simplify?

OpenStudy (anonymous):

no :c

OpenStudy (anonymous):

I can't :c

OpenStudy (melodious):

which grade chemistry ur talking about?

OpenStudy (melodious):

pls say

OpenStudy (joannablackwelder):